v(x,t)= -30cos(30x-15t). Find v when x=1.5 and t=15. Please teach me how to find cos in RADIANT!
Answer #1

v(x,t)= -30cos(30x-15t)

when x=1.5 and t=15

v(1.5,15)= -30cos(30\times1.5-15\times15)=-30cos(-180)

first off change radian into degree

180 rad=180\times180/\pi degree=10313.24 degree

v(1.5,15)= -30cos(30\times1.5-15\times15)=-30cos(-180)=-30cos(-10313.240)=17.95 m/s